OpenStudy (anonymous):

A tank is drained by two pipes. One pipe can empty the tank in 45 mins. while the other can empty in an hour. If the tank is 3/4 filled and both pipes are open , at what time will the tank be emptied?

OpenStudy (anonymous):

pls help :) @jojo4eva

OpenStudy (anonymous):

pls help @mathmath333

OpenStudy (mathmath333):

when the tank is full \(\large\tt \color{black}{\dfrac{x}{45}+\dfrac{x}{60}=1}\)

OpenStudy (mathmath333):

is the answer given in ur book?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

doh i mean \[{\dfrac{x}{45}+\dfrac{x}{60}=\frac{3}{4}}\]

OpenStudy (mathmath333):

well i think here u have to find the value of x from the equation and multiply it by 3/4 not sure lol

OpenStudy (anonymous):

you could be right,

OpenStudy (mathmath333):

\(\large\tt \color{black}{\ddot\smile}\)

OpenStudy (anonymous):

their combined rate is \[\frac{1}{60}+\frac{1}{45}\] and you want to solve \[(\frac{1}{60}+\frac{1}{45})t=\frac{3}{4}\] for \(t\)

OpenStudy (anonymous):

\[\frac{64+45}{60\times 45}t=\frac{3}{4}\] \[t=\frac{3}{4}\times\left(\frac{60\times 45}{60+45}\right)\]